OpenStudy (anonymous):

kiley earned 230 for 40 hours of work. brittany earned $192 for 32 hours. are these pay rates proportional?? explain reasoning PLEASE HELP.

OpenStudy (anonymous):

230/40 = 5.75 192/32= 6.00 so no they are not proportional rates the

OpenStudy (anonymous):

no 6*32=192 so 6*40=230 but that's not true

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

uw

hero (hero):

\[23 \ne 24\]